OpenStudy (anonymous):

Convert each equation to Vertex Formula and Graph each equation showing its vertex. x2 − 5x + 6=0

OpenStudy (anonymous):

vertex form -> y = a(x-h)^2 + k , (h,k) is the vertex point the standard form -> x^2 + bx +c = 0 x2 − 5x + 6=0 (x-3)(x-2) = 0 the graph intersect x-axis at x=3 and x=2 the value of h = 3+2/2 = 5/2 therefore k = (5/2)^2 -5(5/2) + 6 = 25/4 - 25/2 + 6 = -1/4 so now that we know the value of h and k y = a(x - 5/2)^2 - 1/4 to find the value of a, just plug the value of point that's passed by the graph. i'll plug x=1, from the standard form, i'll get y=1-5+6 = 2 then i plug that point (1,2) 2=a(1-5/2)^2 - 1/4 2=a(-3/2)^2 -1/4 2=a(9/4)-1/4 a(9/4) = 9/4 -> a=1 finally we have the complete vertex form y=1(x-5/2)^2 - 1/4 -> y=(x - 5/2)^2 - 1/4

OpenStudy (anonymous):

the graph will be looked like this